OpenStudy (anonymous):

Solve by the elimination method 5x+2y=-7 7x-3y=17

jhonyy9 (jhonyy9):

5x+2y=-7 7x-3y=17 --------- 12x-y=10 --- y=12x-10 -- hence 5x+2(12x-10)=-7 --- 5x+24x-20=-7 --- 29x=13 ---x=13/29 y=12x-10=12(13/29) -10=156/29 -10 =156-290)/29 =-134/29 x=13/29 y=-134/29